OpenStudy (anonymous):

A can of cooldrink holds 330ml. what are dimensions of a can that will hold this volume but which has minimum surface area?

OpenStudy (amistre64):

have you ever noticed that the derivative of Volume IS surface area?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

regardless, define 2 equations, one for volume and one for surface area of a cylinder, take the derivative of surface area and zero it out

OpenStudy (anonymous):

derivate with respect to what?

OpenStudy (amistre64):

you can do implicit since Volume is dependant on h,r

OpenStudy (amistre64):

or, you can use volume to reduce the equation to 1 variable

OpenStudy (amistre64):

\[330=\pi~r^2h\] \[(330)'=\underbrace{2\pi rh~ r'}_{S_a~wrt.r}+\pi r^2h'\] \[S_a=2\pi rh+2\pi r^2\] \[(S_a)'=2\pi r'h+2\pi rh'+4\pi rr'\] \[0=2\pi (r'h+ rh'+2rr')\] \[0=r'(h+2r)+ rh'\]

OpenStudy (amistre64):

on second thought ... i was confusing the sphere and cylindar ... the derivative of a spheres volume is surface area .... too early

OpenStudy (amistre64):

\[0=r'(h+2r)+ rh'\] \[-rh'=r'(h+2r)\] \[-\frac{h'}{r'}=\frac{h+2r}{r}\] from the Volume stuff we can prolly find h'/r' \[0=2\pi rh~ r'+\pi r^2h'\] \[-\pi r^2h'=2\pi rh~ r'\] \[-\frac{h'}{r'}=2\pi \frac{rh}{\pi r^2}\] \[-\frac{h'}{r'}=2\frac{h}{r}\]

OpenStudy (amistre64):

so, 2h = 2r+h, when h=2r

OpenStudy (amistre64):

use this to solve for r:\[330=\pi~r^2h\] use this to solve for r:\[330=2\pi~r^3\]

OpenStudy (amistre64):

another view is that Sa is a parabola \[S_a=2\pi rh+2\pi r^2\] when \(r=\Large \frac{-2\pi h}{2(2\pi)}\) we are at a minimum: r = -h/2, to which we get again that |h| = 2r 330 = pi (-h/2)^2 h 330 = pi h^3/4 h = cbrt(4(330)/pi)